OpenStudy (anonymous):

Salina purchased a dining room set for $1,730 using a 12-month deferred payment plan with an interest rate of 23.45%. She did not make any payments during the deferment period. What will the total cost of the dining room set be if she must pay off the dining room set within two years after the deferment period?

OpenStudy (anonymous):

@kitty828 Do you know which formula to apply?

OpenStudy (anonymous):

I have no idea how to do this problem

OpenStudy (anonymous):

I ask about the formula you've learned from your class, not how to solve it!

OpenStudy (anonymous):

M=B i(1+i)nt over (1+i)nt-1 B=P(1+i)nt

OpenStudy (anonymous):

I'm not sure I couldn't figure that out /: I couldn't figure out the i or the nt either

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

As the rule of thumb, 23.45% is annual rate -> How can you find monthly rate?

OpenStudy (anonymous):

I am very confused. these are the options for answers $2,754.72 $4,132.08 $1,730.00 $2,182.26

OpenStudy (anonymous):

No, I guide you to think! What the relation between year and month?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

You probably know how many months per year :)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Now, you know 23.45% is the interest rate per year Could you find interest rate per month = ...?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

1.9

OpenStudy (anonymous):

Pretty close, except you forget to divide by 100 :)

OpenStudy (anonymous):

.0195?

OpenStudy (anonymous):

Yup, you're making a progress :) Can you find the time now ?

OpenStudy (anonymous):

the time? how would I do that?

OpenStudy (anonymous):

Convert year into months, because the interest based on months :)

OpenStudy (anonymous):

so 12?

OpenStudy (anonymous):

Check for the period ?

OpenStudy (anonymous):

sorry 24. i didnt relize it was 2 years

OpenStudy (anonymous):

Sorry I'm on the phone solving some family trouble!

OpenStudy (anonymous):

its ok thank you for helping me!

OpenStudy (anonymous):

So now you know i and nt, correct!

OpenStudy (anonymous):

so this is the correct answer? $4,132.08

OpenStudy (anonymous):

Can you plug into the formula: M = B i(1+i)^nt / (1+i)^(nt-1)

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Could you show your calculation?

OpenStudy (anonymous):

I don't know i'm so confused at this point. Where do I need to plug that number into on the formula? into i?

OpenStudy (anonymous):

1. Balance from 12-month deferred payment plan: B = P (1 + i) ^nt = 1,730 ( 1+ .0195 ) ^ 12 = $2,182.26 2. Monthly payment after 12 months: M = B i (1 + i ) ^nt / (1+i)^ nt - 1 = $2,182.26 * .0195 ( 1.0195) ^24 / [ ( 1.0195) ^24 -1 ] = $114.78 3. Total cost: T = 24 * $114.78 = $ 2, 754.72

OpenStudy (anonymous):

this answer is correct